hallm:

The formula for the perimeter of a rectangle is P=2(l+w). Solve for l. l=P2+w l=p−w2 l=p2−w l=2P−w

hallm:

@shadow please help

jhonyy9:

P=2(l+w) distribute P=2l +2w subtract from both sides 2w P -2w = 2l divide both sides by 2 (P -2w)/2 =l hope helped